xx ConverttoHydroBase: A Module for Converting IllinoisGRMHD variables into HydroBase variables, for Analysis

ConvertxtoxHydroBase: A Module for Converting IllinoisGRMHD variables into HydroBase variables, for Analysis

Original author: Zachariah B. Etienne.

Date: 2015-10-12 12:00:00 -0600 (Mon, 12 Oct 2015)

Abstract

IllinoisGRMHD and HydroBase variables are incompatible; The former uses 3-velocity defined as , and the latter uses the Valencia formalism definition of .

Define the Valencia formalism’s definition of to be ””, and IllinoisGRMHD’s definition ”” Then

which comes from Eq 11 in the IllinoisGRMHD code announcement paper:

http://arxiv.org/pdf/1501.07276.pdf.

Similarly,

In addition, IllinoisGRMHD needs the A-fields to be defined on staggered grids, and HydroBase does not yet support this option. The staggerings are defined in Table 1 of the IllinoisGRMHD code announcement paper:

http://arxiv.org/pdf/1501.07276.pdf (page 15).

The long-term goal should be to adjust HydroBase and IllinoisGRMHD so that this thorn is no longer necessary.

1 Parameters

11convertxtoxhydrobasexevery 10Scope: privatex01  INT

11Description: How often to convert IllinoisGRMHD primitive variables to HydroBase (Valencia formulation) primitive variables? Needed for some ET-based diagnostics. NOT needed for pure IllinoisGRMHD runs.

11  Range  10  x01  Default: (none)
110:*  10zero (disable) or positive (every N iterations)

11

2 Interfaces

General Implements: convertxtoxhydrobase

3 Schedule This section lists all the variables which are assigned storage by thorn WVUThorns/ConvertxtoxHydroBase. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage NONE